# Probing AGN Feedback in Dwarf Galaxies with Spatially Resolved NIR Coronal Lines from JWST

### v1

@@ 5.1 Non-Uniform Frequency of the Coronal Lines
Previous studies of CLs in dwarf galaxies have found that few systems display all CLs detectable in a given spectral window simultaneously (2006A&A...457...61R). The most frequent CLs detections are from Si and S, with the lower-ionization lines (I.P. 300 eV) typically being stronger. Lines from other elements such as Ca, Al, and Fe are generally weaker and less often detected. Ca and Al can be suppressed by metallicity and depletion onto dust, while Fe may be less affected. Other CLs, such as those from Mg, have rarely been studied because their wavelengths were largely inaccessible in the pre-JWST era. In massive galaxies, it was initially assumed that the frequency and strength of CLs should depend on the orientation of the torus with respect to the line of sight (1998ApJ...497L...9M; 2000AJ....119.2605N; 2009MNRAS.397..172G), but subsequent work has found little to no difference in the number of detected CLs between Type 1 and Type 2 AGN (2006A&A...457...61R; 2010MNRAS.405.1315M; 2011ApJ...743..100R). The absence of CLs above certain ionization energies may instead point to a softer ionizing continuum. In some cases, higher-I.P. lines are detected while lower-I.P. lines of different species are absent (2017MNRAS.467..540L; 2024ApJ...975...60A), which can be explained by differences in elemental abundances or lack of sensitivity from ground-based NIR instruments. In our sample, we mostly find that, within a given species, there are very few cases where a higher-I.P. line is detected but the corresponding lower-I.P. line is not.
@@
To investigate the origin of the non-uniformity in CL detections, we first examine diagnostics of the ionizing source. J0842 is the only target that falls in the composite region of the BPT diagram when using Keck/LRIS fluxes from 2019ApJ...884...54M. It may either lack a powerful, hard-spectrum AGN relative to the others, or its AGN signatures may be significantly diluted by stellar contamination. All other targets (J0906, J0954, J1009) lie in the Seyfert region of the BPT diagram (2019ApJ...884...54M). Both J0906 and J0954 have strong X-ray detections consistent with AGN activity (2017ApJ...836...20B), and J0906 additionally hosts a parsec-scale radio jet (2020MNRAS.495L..71Y). J0906 also has the most luminous CLs in our sample, and the median FWHM of its narrow CL component exceeds the median FWHM of both the narrow and broad components in the other three targets, consistent with a powerful central engine and energetic nuclear gas.
@@
J1009 exhibits the richest coronal-line spectrum, with 16 lines detected. As shown in Section 4.2, its log values are less negative than those of J0954, likely contributing to the larger number of CLs, including [S XI] (I.P. = 447 eV), which is not detected in J0954. J1009 also has the lowest stellar mass among our targets and, based on black hole–stellar mass relations (2015ApJ...813...82R), is expected to host a relatively low-mass black hole. Previous studies have shown that lower-mass black holes are more likely to exhibit multiple strong high-ionization CLs (2018ApJ...861..142C). As black hole mass decreases, the Schwarzschild radius decreases and the accretion disk temperature increases, shifting the ionizing continuum to higher energies and enhancing the production of high-ionization lines. The CL detections in J1009 are consistent with this scenario.
@@
We next consider the role of dust. Balmer decrements (assuming intrinsic ratios suitable for AGN) suggest that these galaxies should not have high dust extinction (2020ApJ...905..166L; 2021ApJ...911...70B). However, the E(B–V) values inferred from the Pa/Pa ratio (Table 3) for J0906 and J0954 are significantly higher; these were derived assuming an intrinsic Pa/Pa ratio of 2.133, more appropriate for H II regions. None of the four targets show strong large-scale spatial variations between spaxels in their extinction maps.
@@
5.2 Coronal line region in dwarf galaxies
Determining the size and physical nature of the region where these CLs are formed (the coronal line region; CLR) has long been a challenge. Early work placed the CLR in a transitional zone between the broad-line region (BLR) and narrow-line region (NLR) (1981MNRAS.195..787P; 1984MNRAS.208..347P; 1988ApJS...67..373E), but subsequent observations showed that CLs in massive galaxies can extend to hundreds of parsecs, implying that the CLR can span a substantial fraction of the NLR (2005MNRAS.364L..28P; 2006ApJ...653.1098R; 2011ApJ...739...69M; 2021ApJ...920...62N).
@@ 5.2 Coronal line region in dwarf galaxies
We use Kendall’s (appropriate for our small sample) to test for correlations between ionization potential (I.P.) and the extents (determined using the method listed in Section 3). The element-wise results are weakly consistent with smaller extents for higher-ionization lines, but are not statistically significant (likely due to the small sample size). Si shows the strongest possible trend (, ), S is also negative (, ), and Mg and the remaining species show negative , while Ca shows no trend (, ). The Ca lines despite their relatively low I.P. compared to many other CLs, have among the smallest extents. This might possibly be linked to the strong depletion of Ca onto dust grains at larger radii (2004ApJS..153...75G; 2024ApJ...976..130M; 2025ApJ...984..170M), which reduces the gas-phase Ca available for ionization. Pooling all lines in a stratified Kendall’s that preserves galaxy identity yields (, 95% CI [, ]), which is consistent with the same trend but not significant, which may be due to the fact that mixing different elements introduces abundance and depletion effects. Overall, the statistics point toward decreasing spatial extent with increasing I.P., but a larger sample is required for a definitive constraint.
@@
If these weak trends turn out to be real, they might be interpreted in the context of photoionization theory and the density structure of gas around AGN (1997ApJ...487..122F; 2011ApJ...743..100R; 2013MNRAS.430.2411M). In ionization-bounded conditions, the CLR size is limited by the available ionizing photons: high-I.P. CLs form closer to the nucleus, while lower-I.P. CLs can be sustained at larger radii. In gas-density-bound conditions, ionizing photons are plentiful but the gas density is too low for efficient collisional excitation into the upper levels, linking the spatial distribution of a line to its critical density . For , line emission scales roughly with , so CLs with higher are preferentially emitted in denser inner regions and lower- lines can extend farther out.
@@
We find no significant correlation between CL extent and . This suggests that the CL extents might be primarily governed by the ionizing radiation field rather than by gas-density limitations. Combined with the results of Section 4.2, this supports a picture in which AGN photoionization dominates the production and radial extent of CLs.
@@
The average CL extent in our sample is kpc (Figure 9). For J0954, the outflow radius of kpc inferred from blueshifted UV absorption (C II, C IV, Si II, Si IV) by 2024ApJ...965..152L is consistent with our coronal-line extents, indicating that the CLs might trace the ionized outflow. When multiple components are present, the broader component has typical extents of kpc or less. We can compare our CL extents with measurements in massive galaxies, assuming those have similar sensitivity to our measurements (2011ApJ...739...69M; 2005MNRAS.364L..28P; 2006ApJ...653.1098R). The absolute CL extents in our dwarfs are similar, but because the hosts are smaller, the CLR appears to occupy a larger fraction of the galaxy (10% versus ). However, optical CLs in massive galaxies have been shown to be more extended (0.7–3 kpc; 2017MNRAS.470.2845R; 2021ApJ...920...62N; 2025MNRAS.538.2800R) so our NIR-based comparison is conservative.
5.3 Correlation with ionized-gas outflows
@@
Although the AGN accretion disk is expected to generate abundant photons capable of producing CLs, recent systematic surveys investigating optical CLs in large samples of nearby galaxies have found such emission to be exceedingly rare within the general galaxy population. This rarity persists even among galaxies identified as AGNs using standard narrow-line ratio diagnostics, at current survey sensitivities (2021ApJ...922..155M; 2021ApJ...920...62N; 2022ApJ...936..140R; 2023ApJ...945..127N). Furthermore, these surveys indicate that CL emission preferentially appears in galaxies with low dust extinction (2004ApJS..153...75G; 2025ApJ...984..170M), suggesting that dust may contribute to suppressing CL emission.
@@
Recent photoionization models by 2024ApJ...976..130M show that gas-phase depletion of refractory elements onto dust grains, along with dust absorption of ionizing photons, can significantly impact CL emission, reducing line luminosities by up to three orders of magnitude compared to dust-free conditions. Consequently, strong CL emission likely occurs only in environments where sufficient grain destruction takes place within highly ionized gas (2003AJ....125.1729N). All dwarf galaxies with CL in this sample also exhibit low levels of dust extinction in the central regions, at least based on their Balmer decrements (Table 3).
@@
All galaxies in this work show some of the fastest ionized-gas outflows reported in dwarf galaxies, as traced by their broad, blue-shifted [O III] emission-line profiles (2019ApJ...884...54M; 2020ApJ...905..166L). The [O III] outflow component contributes 5–30% of the total [O III] flux in these galaxies (2019ApJ...884...54M). Both J0906 and J0954 require three Gaussian components to model the [O III] emission lines, revealing a broad component with an average FWHM of 800 km s-1. J0906 and J0954 have maximum outflow velocities of 150 and 80 km s-1, respectively, where the negative sign indicates blue-shifted velocity centroids relative to the systemic velocity. J1009 has a maximum outflow velocity of 60 km s-1 (see also Section 2.1 and Table 1 for more details on the ionized gas outflow properties). The warm ionized-gas mass outflow rates range from – M⊙ yr-1, and the kinetic energy outflow rates vary between to erg s-1, which are comparable to those of more luminous AGN in massive systems when normalized by AGN bolometric luminosities.
@@
Outflows could enhance CL emission by clearing out dust from the interior of the galaxies. This was previously suggested in a study of bulgeless galaxies with AGN, where 2022ApJ...931...69B found that all the galaxies that have AGN-driven outflows also show CL emission, suggesting a correlation between the two. In a systematic study of galaxies with ionized-gas outflows, 2025ApJ...984..170M showed that galaxies with CL emission exhibit a significantly higher incidence and luminosity of ionized outflows, traced by [O III] 5007, compared to a control sample of non-CL-emitting galaxies. They also found that there exists strong correlations between CL properties (luminosity and FWHM) and outflow velocity. They find that the CL emitters have systematically lower intrinsic extinction toward the ionized gas compared to the controls, supporting a scenario in which dust destruction in AGN-driven outflows facilitates efficient CL emission. Although there is a possibility that this correlation is a by-product of AGN power, 2025ApJ...984..170M do not find a correlation between the total AGN bolometric luminosity and the CL luminosity (which we also find in this work.)
@@
In our analysis, we also find several correlations between infrared CL properties and the characteristics of ionized gas outflows traced by [O III] (2020ApJ...905..166L) in our sample galaxies. The detected CLs are spatially extended along the same direction as the outflows detected in [O III] (Figure 11). The extended CLs also exhibit a largely biconical morphology.
@@
Furthermore, we apply a Kendall’s correlation test to determine if a significant correlation exists between the various quantities, such as the CL luminosity and CL FWHM and ionized gas outflow luminosity, [O III] FWHM and the [O III] outflow energetics. The ionized gas outflow luminosity is calculated from the outflow masses given in Table 4 of 2020ApJ...905..166L using Equation 7 in the same work. The CL luminosities and FWHM are calculated from the integrated spectra using a 0.3″central aperture for all the detected CL (we show a subsample of the different CL in Figures 12 and 13, but the correlations are present across all the CLs). We find the strongest correlations (median Kendall’s coefficient = 0.42 and p-value = 0.05) between the CL luminosity and the AGN outflow luminosity (Figure 12) as well as between the CL luminosity and the most energetic outflows, given as the kinetic energy injection rate (dE/dt) normalized by the AGN bolometric luminosity (Figure 13). We do not find a significant correlation between the total AGN bolometric and [O III] luminosities and the CL luminosities. We also find a significant correlation (median Kendall’s coefficient 0.4 and p-value 0.05) between CL luminosity and the FWHM of the [O III] outflow component. We do not find a significant correlation between the FWHM of CL and that of the [O III] outflow component, which agrees with the results from 2025ApJ...984..170M.
@@
While some previous studies have suggested that CLs with higher I.P. tend to exhibit broader profiles (2002ApJ...579..214R; 2023ApJ...942L..37A), we do not find a statistically significant correlation between either FWHM or W80 and the I.P. or the critical density (ncrit) of the lines (Figure 14). However, we do find that targets with larger W80 values in the broad [O III] component also show correspondingly large W80 values across their CLs (Figure 14).
@@
5.4 Outflows traced by Coronal Lines
We find that several CLs in the observed dwarf galaxies have a second, broad component suggestive of possible outflow. While CL broadening may reflect AGN activity, the presence of outflows detected by the [O III] lines necessitates investigating whether even the highest-ionization gas participates in the outflow. The velocity offsets of the broad CL range from 150 to 150 km s-1, with W80 values ranging from 300 - 500 km s-1. We find that, overall, the broad CL components have smaller velocities as compared to lines such as [S III] (0.9533 m) and He II (1.0125 m), which have similar I.P.s and ne to [O III]. A more detailed analysis of outflows in other lines will be presented in a future study.

### v2

@@ V.1 Non-Uniform Frequency of the Coronal Lines
Previous studies of CLs in dwarf galaxies have found that few systems display all CLs detectable in a given spectral window simultaneously (Riffel et al., 2006). The most frequent CLs detections are from Si and S, with the lower-ionization lines (I.P. 300 eV) typically being stronger. Lines from other elements such as Ca, Al, and Fe are generally weaker and less often detected. Ca and Al can be suppressed by metallicity and depletion onto dust, while Fe may be less affected. Other CLs, such as those from Mg, have rarely been studied because their wavelengths were largely inaccessible in the pre-JWST era. In massive galaxies, it was initially assumed that the frequency and strength of CLs should depend on the orientation of the torus with respect to the line of sight (Murayama and Taniguchi, 1998; Nagao et al., 2000; Gelbord et al., 2009), but subsequent work has found little to no difference in the number of detected CLs between Type 1 and Type 2 AGN (Riffel et al., 2006; Mazzalay et al., 2010; Rodríguez-Ardila et al., 2011). The absence of CLs above certain ionization energies may instead point to a softer ionizing continuum. In some cases, higher-I.P. lines are detected while lower-I.P. lines of different species are absent (Lamperti et al., 2017; Aravindan et al., 2024), which can be explained by differences in elemental abundances or lack of sensitivity from ground-based NIR instruments. In our sample, we mostly find that, within a given species, there are very few cases where a higher-I.P. line is detected but the corresponding lower-I.P. line is not.
@@
To investigate the origin of the non-uniformity in CL detections, we first examine diagnostics of the ionizing source. J0842 is the only target that falls in the composite region of the BPT diagram when using Keck/LRIS fluxes from Manzano-King et al. 2019. It may either lack a powerful, hard-spectrum AGN relative to the others, or its AGN signatures may be significantly diluted by stellar contamination. All other targets (J0906, J0954, J1009) lie in the Seyfert region of the BPT diagram (Manzano-King et al., 2019). Both J0906 and J0954 have strong X-ray detections consistent with AGN activity (Baldassare et al., 2017), and J0906 additionally hosts a parsec-scale radio jet (Yang et al., 2020). J0906 also has the most luminous CLs in our sample, and the median FWHM of its narrow CL component exceeds the median FWHM of both the narrow and broad components in the other three targets, consistent with a powerful central engine and energetic nuclear gas.
@@
J1009 exhibits the richest coronal-line spectrum, with 16 lines detected. As shown in Section IV.2, its log values are less negative than those of J0954, likely contributing to the larger number of CLs, including [S XI] (I.P. = 447 eV), which is not detected in J0954. J1009 also has the lowest stellar mass among our targets and, based on black hole–stellar mass relations (Reines and Volonteri, 2015), is expected to host a relatively low-mass black hole. Previous studies have shown that lower-mass black holes are more likely to exhibit multiple strong high-ionization CLs (Cann et al., 2018). As black hole mass decreases, the Schwarzschild radius decreases and the accretion disk temperature increases, shifting the ionizing continuum to higher energies and enhancing the production of high-ionization lines. The CL detections in J1009 are consistent with this scenario.
@@
We next consider the role of dust. Balmer decrements (assuming intrinsic ratios suitable for AGN) suggest that these galaxies should not have high dust extinction (Liu et al., 2020; Bohn et al., 2021). However, the E(B–V) values inferred from the Pa/Pa ratio (Table 3) for J0906 and J0954 are significantly higher; these were derived assuming an intrinsic Pa/Pa ratio of 2.133, more appropriate for H II regions. None of the four targets show strong large-scale spatial variations between spaxels in their extinction maps.
@@
V.2 Coronal line region in dwarf galaxies
Determining the size and physical nature of the region where these CLs are formed (the coronal line region; CLR) has long been a challenge. Early work placed the CLR in a transitional zone between the broad-line region (BLR) and narrow-line region (NLR) (Pelat et al., 1981; Penston et al., 1984; Evans, 1988), but subsequent observations showed that CLs in massive galaxies can extend to hundreds of parsecs, implying that the CLR can span a substantial fraction of the NLR (Prieto et al., 2005; Rodríguez-Ardila et al., 2006; Müller-Sánchez et al., 2011; Negus et al., 2021).
@@ V.2 Coronal line region in dwarf galaxies
We use Kendall’s (appropriate for our small sample) to test for correlations between ionization potential (I.P.) and the extents (determined using the method listed in Section III). The element-wise results are weakly consistent with smaller extents for higher-ionization lines, but are not statistically significant (likely due to the small sample size). Si shows the strongest possible trend (, ), S is also negative (, ), and Mg and the remaining species show negative , while Ca shows no trend (, ). The Ca lines despite their relatively low I.P. compared to many other CLs, have among the smallest extents. This might possibly be linked to the strong depletion of Ca onto dust grains at larger radii (Groves et al., 2004b; McKaig et al., 2024; Matzko et al., 2025), which reduces the gas-phase Ca available for ionization. Pooling all lines in a stratified Kendall’s that preserves galaxy identity yields (, 95% CI [, ]), which is consistent with the same trend but not significant, which may be due to the fact that mixing different elements introduces abundance and depletion effects. Overall, the statistics point toward decreasing spatial extent with increasing I.P., but a larger sample is required for a definitive constraint.
@@
If these weak trends turn out to be real, they might be interpreted in the context of photoionization theory and the density structure of gas around AGN (Ferguson et al., 1997; Rodríguez-Ardila et al., 2011; Mazzalay et al., 2013). In ionization-bounded conditions, the CLR size is limited by the available ionizing photons: high-I.P. CLs form closer to the nucleus, while lower-I.P. CLs can be sustained at larger radii. In gas-density-bound conditions, ionizing photons are plentiful but the gas density is too low for efficient collisional excitation into the upper levels, linking the spatial distribution of a line to its critical density . For , line emission scales roughly with , so CLs with higher are preferentially emitted in denser inner regions and lower- lines can extend farther out.
@@
We find no significant correlation between CL extent and . This suggests that the CL extents might be primarily governed by the ionizing radiation field rather than by gas-density limitations. Combined with the results of Section IV.2, this supports a picture in which AGN photoionization dominates the production and radial extent of CLs.
@@
The average CL extent in our sample is kpc (Figure 9). For J0954, the outflow radius of kpc inferred from blue-shifted UV absorption (C II, C IV, Si II, Si IV) by Liu et al. 2024 is consistent with our coronal-line extents, indicating that the CLs might trace the ionized outflow. When multiple components are present, the broader component has typical extents of kpc or less. We can compare our CL extents with measurements in massive galaxies, assuming those have similar sensitivity to our measurements (Müller-Sánchez et al., 2011; Prieto et al., 2005; Rodríguez-Ardila et al., 2006). The absolute CL extents in our dwarfs are similar, but because the hosts are smaller, the CLR appears to occupy a larger fraction of the galaxy (10% versus ). However, optical CLs in massive galaxies have been shown to be more extended (0.7–3 kpc; Rodríguez-Ardila et al. 2017; Negus et al. 2021; Rodríguez-Ardila et al. 2025) so our NIR-based comparison is conservative.
V.3 Correlation with ionized-gas outflows
@@
Although the AGN accretion disk is expected to generate abundant photons capable of producing CLs, recent systematic surveys investigating optical CLs in large samples of nearby galaxies have found such emission to be exceedingly rare within the general galaxy population. This rarity persists even among galaxies identified as AGNs using standard narrow-line ratio diagnostics, at current survey sensitivities (Molina et al., 2021; Negus et al., 2021; Reefe et al., 2022; Negus et al., 2023). Furthermore, these surveys indicate that CL emission preferentially appears in galaxies with low dust extinction (Groves et al., 2004b; Matzko et al., 2025), suggesting that dust may contribute to suppressing CL emission.
@@
Recent photoionization models by McKaig et al. 2024 show that gas-phase depletion of refractory elements onto dust grains, along with dust absorption of ionizing photons, can significantly impact CL emission, reducing line luminosities by up to three orders of magnitude compared to dust-free conditions. Consequently, strong CL emission likely occurs only in environments where sufficient grain destruction takes place within highly ionized gas (Nagao et al., 2003). All dwarf galaxies with CL in this sample also exhibit low levels of dust extinction in the central regions, at least based on their Balmer decrements (Table 3).
@@
All galaxies in this work show some of the fastest ionized-gas outflows reported in dwarf galaxies, as traced by their broad, blue-shifted [O III] emission-line profiles (Manzano-King et al., 2019; Liu et al., 2020). The [O III] outflow component contributes 5–30% of the total [O III] flux in these galaxies (Manzano-King et al., 2019). Both J0906 and J0954 require three Gaussian components to model the [O III] emission lines, revealing a broad component with an average FWHM of 800 km s-1. J0906 and J0954 have maximum outflow velocities of 150 and 80 km s-1, respectively, where the negative sign indicates blue-shifted velocity centroids relative to the systemic velocity. J1009 has a maximum outflow velocity of 60 km s-1 (see also Section II.1 and Table 1 for more details on the ionized gas outflow properties). The warm ionized-gas mass outflow rates range from – M⊙ yr-1, and the kinetic energy outflow rates vary between to erg s-1, which are comparable to those of more luminous AGN in massive systems when normalized by AGN bolometric luminosities.
@@
Outflows could enhance CL emission by clearing out dust from the interior of the galaxies. This was previously suggested in a study of bulgeless galaxies with AGN, where Bohn et al. 2022 found that all the galaxies that have AGN-driven outflows also show CL emission, suggesting a correlation between the two. In a systematic study of galaxies with ionized-gas outflows, Matzko et al. 2025 showed that galaxies with CL emission exhibit a significantly higher incidence and luminosity of ionized outflows, traced by [O III] 5007, compared to a control sample of non-CL-emitting galaxies. They also found that there exists strong correlations between CL properties (luminosity and FWHM) and outflow velocity. They find that the CL emitters have systematically lower intrinsic extinction toward the ionized gas compared to the controls, supporting a scenario in which dust destruction in AGN-driven outflows facilitates efficient CL emission. Although there is a possibility that this correlation is a by-product of AGN power, Matzko et al. 2025 do not find a correlation between the total AGN bolometric luminosity and the CL luminosity (which we also find in this work.)
@@
In our analysis, we also find several correlations between infrared CL properties and the characteristics of ionized gas outflows traced by [O III] (Liu et al., 2020) in our sample galaxies. The detected CLs are spatially extended along the same direction as the outflows detected in [O III] (Figure 11). The extended CLs also exhibit a largely biconical morphology.
@@
Furthermore, we apply a Kendall’s correlation test to determine if a significant correlation exists between the various quantities, such as the CL luminosity and CL FWHM and ionized gas outflow luminosity, [O III] FWHM and the [O III] outflow energetics. The ionized gas outflow luminosity is calculated from the outflow masses given in Table 4 of Liu et al. 2020 using Equation 7 in the same work. The CL luminosities and FWHM are calculated from the integrated spectra using a 0.3″central aperture for all the detected CL (we show a subsample of the different CL in Figures 12 and 13, but the correlations are present across all the CLs). We find the strongest correlations (median Kendall’s coefficient = 0.42 and p-value = 0.05) between the CL luminosity and the AGN outflow luminosity (Figure 12) as well as between the CL luminosity and the most energetic outflows, given as the kinetic energy injection rate (dE/dt) normalized by the AGN bolometric luminosity (Figure 13). We do not find a significant correlation between the total AGN bolometric and [O III] luminosities and the CL luminosities. We also find a significant correlation (median Kendall’s coefficient 0.4 and p-value 0.05) between CL luminosity and the FWHM of the [O III] outflow component. We do not find a significant correlation between the FWHM of CL and that of the [O III] outflow component, which agrees with the results from Matzko et al. 2025.
@@
While some previous studies have suggested that CLs with higher I.P. tend to exhibit broader profiles (Rodríguez-Ardila et al., 2002; Armus et al., 2023), we do not find a statistically significant correlation between either FWHM or W80 and the I.P. or the critical density (ncrit) of the lines (Figure 14). However, we do find that targets with larger W80 values in the broad [O III] component also show correspondingly large W80 values across their CLs (Figure 14).
@@
V.4 Outflows traced by Coronal Lines
We find that several coronal lines (CLs) in the observed dwarf galaxies require a second, broad component. This may indicate that the coronal-line region (CLR) arises closer to the central AGN than the narrow-line region (NLR), leading to intrinsically broader profiles (Rodríguez-Ardila et al., 2002; Mazzalay et al., 2010). Broad CLs have also been reported in tidal disruption events (TDEs; Wang et al. 2011), but the lack of fading between the Keck/NIRES observations taken in 2019 and JWST/NIRSpec observations taken in 2024 (see Section IV.1) makes a TDE origin unlikely in our case. In massive AGN, CLs are often associated with outflows, and their typically blue-shifted, broadened profiles are frequently interpreted as signatures of outflowing gas (Müller-Sánchez et al., 2011; Mazzalay et al., 2013; Rodríguez-Ardila et al., 2017; May et al., 2018; Fonseca-Faria et al., 2021; Speranza et al., 2022).
Bohn et al. (2021) detected a secondary broad component in the [Si VI] emission lines of J0906, J0954, and J1009 using Keck/NIRES. They measured velocity offsets up to 600 km s-1 and maximum W80 values of 1000 km s-1 for the broad component, strongly suggestive of outflowing coronal gas. We do not recover comparably high velocities in [Si VI] with JWST/NIRSpec. However, individual spaxels in the NIRSpec IFU cubes are affected by “wiggles” (Section II.2), which complicates separating weak, broad wings from spectral artifacts and likely leads us to underestimate the outflow velocities. Nevertheless, the presence of a broad component in both the CLs and the [O III] lines in the ground-based data motivates a closer examination of whether even the highest-ionization gas participates in the outflow.
The broad CL components in this work show velocity offsets of 150 to 150 km s-1 and W80 values of 300–500 km s-1. These values should be treated as lower limits, since the outflow wings are difficult to disentangle from the “wiggles.” Overall, the broad CL components have smaller velocities than lines such as [S III] 0.9533 m and He II 1.0125 m, which have similar ionization potentials and electron densities to [O III]. A more detailed outflow analysis using additional emission lines will be presented in a future study.
The triangles are measurements from [O III] (Liu et al., 2020). The squares with crosses indicate energetics calculated assuming a filling factor using equations from Müller-Sánchez et al. 2011, also used in Rodríguez-Ardila et al. 2017. Across the sample, coronal-line–based outflow energetics estimated using the filling factor (squares with crosses) approximation are 1.0–2.0 dex higher than those estimated using the [Si VI] luminosities (filled squares).